OpenStudy (anonymous):

a $16000 car was decreased 15% each year how much is it after 6 years

OpenStudy (btaylor):

\[A = P(1-r)^t \rightarrow A = 16000(1-.15)^6 \rightarrow A = 16000(0.85)^6 \rightarrow A = 16000(0.377)=$6034.39\]

OpenStudy (anonymous):

what does it equal the box showing your answer is cut off

OpenStudy (btaylor):

$6034.39 sorry about that!